Jamila and Tara go out for a bicycle ride. They leave from the same place riding inopposite directions. Tara is riding 10 miles per hour faster than Jamila. After riding for 6 hours, they are 180 miles apart from each other. At what speed was Jamila riding?
Answer by Paul(988)

Tara's speed = x+10
Jamila's speed = x
Equation:
6(x+x+10)=180
12x+60=180
12x=120
x=10
10+10=20
Hence, one is travelling 10mph and one is traveling 20mph.
Paul.